The table shows the number of years of education of parent and the number of years of education of their child.  

Years of Education of Parent

13

9

7

12

12

10

11

Years of Education of their Child

13

11

7

16

17

9

17

Determine the correlation coefficient and briefly describe the relationship between the two variables.
Conduct a hypothesis test to determine if there is a significant linear correlation between the two variables. What is the critical value?
Conduct a hypothesis test to determine if there is a significant linear correlation between the two variables. What is the test statistic?
Conduct a hypothesis test to determine if there is a significant linear correlation between the two variables. What is the p-value?
If there is a significant linear correlation between the variables, determine the regression equation
Answer #1

Correlation Coefficient Calculation

Step 1: Find X Y, X and Y2 as it was done in the table below. X-Y X X- Y -Y 144 256 16 17 192 204 12 Step 2: Find the sum of

r = 0.7729

Critical value

The sample size is n = 7, so then the number of degrees of freedom is df = n - 2 = 7 - 2 = 5

α = 0.05

Critical t value for df = 5 and α = 0.05 is +/- 2.5706 (Obtained using calculator. Screenshot attached)

Degrees of freedom: 5 Probability level: 0.05 Calculate! t-value (right-tail): 2.01504837 t-value (two-tailed): +/- 2.5705818

Test statistic

The corresponding t-statistic to test for the significance of the correlation is: 7-2 t=rvm-3 = 0.7200 10 72g = 2.724 V1-0.77

Test statistic = 2.724

P-value

P-value corresponding to t = 2.724 and df = 5 is 0.0416 (Obtained using calculator)

Since p-value = 0.0416 < α = 0.05, we reject null hypothesis. Thus, there is a significant linear correlation between the variables

Regression equation

We will find an equation of the regression line in 4 steps. Step 1: Find X Y and X2 as it was done in the table below. Y X-Y

The equation of the regression line is:

y = -3 + 1.5*x